# A Survey of Cross-domain Graph Learning: Progress and Future Directions

### v1

@@ VI Future Directions
VI-A Feature Alignment.
The correlation between feature dimensions and semantics across graph data from different domains is a crucial aspect to consider in CDGL. Misaligned feature dimensions will hinder the training of a general graph model capable of handling graph data from various domains, while misaligned semantics will prevent the trained graph model from capturing shared characteristics across domains, thereby limiting the performance of CDGL.
In existing CDGL methods, those targeting limited-scale scenarios can often skip this issue, as the graph data from source and target domains typically share identical feature dimensions by nature, such as in molecular networks [17, 36]. For conditional scenarios, which mainly focus on graph data in recommendation systems [70] and text-attributed graphs (TAGs) [38, 41], feature alignment challenges are circumvented in different ways: Graph data in recommendation systems may come from different domains, but they are all built on the basis of users and items, thus avoiding feature alignment problems. For TAGs, features can be easily aligned using embeddings generated by large language models (LLMs).
However, these methods are insufficient for addressing open CDGL, where feature dimensions and semantics are highly complex and heterogeneous across domains. Although some methods have begun exploring feature alignment solutions for more open scenarios, the research is still in its early stages. For instance, methods like ProG [93], MDGPT [72], and CDFS-GAD [40] have attempted to align feature dimensions across different datasets using techniques such as SVD or MLP. However, even after aligning feature dimensions, the aligned features may still lack semantic equivalence. Therefore, proposing a universal feature alignment approach to handle datasets from different domains is a key research direction towards achieving effective open CDGL.
VI-B Dealing with the Scale of Cross-domain Datasets.
@@
VI-C Overcoming data leakage.
Table I demonstrates that LLMs are receiving increasing attention in CDGL, exploring their potential capabilities in this area. However, the introduction of LLMs inevitably brings about the issue of data leakage, which has become a focal point of discussion [94]. Specifically, since LLMs are pre-trained on extensive text corpora, it’s likely that they may have seen and memorized at least part of the test data from common benchmark datasets, especially in citation networks. This undermines the reliability of current studies that rely on earlier benchmark datasets. Moreover, [95] demonstrates that specific prompts could potentially enhance the ”activation” of LLMs’ corresponding memory, thereby influencing evaluation outcomes. Both [96] and [45] have attempted to avoid the data leakage issue by collecting a new citation dataset, ensuring that the test papers are sampled from time periods after ChatGPT’s data cut-off. However, these efforts remain limited to the citation domain, and the impact of graph structures in their datasets is not significant. Therefore, it’s crucial to reconsider the methods employed to accurately evaluate the performance of LLMs on graph-related tasks. A fair, systematic, and comprehensive benchmark is also needed.
VI-D Improving interpretability.
Interpretability, also known as explainability, refers to the ability to explain or present a model’s behavior in terms that humans can understand [91]. In CDGL, most methods utilize Graph Neural Networks (GNNs). However, GNNs are typically treated as black boxes [97] and lack interpretability. While these methods consider the relationships among various pieces of information in the graph during their design and have experimentally demonstrated effectiveness in CDGL, their improvements are often based on intuition and a high-level understanding of graph data. Consequently, their interpretability remains poor. To enhance interpretability, CDGL approaches that integrate Large Language Models (LLMs) have been proposed. This is primarily due to LLMs’ reasoning and explanatory abilities, which enable them to produce user-friendly explanations for graph reasoning tasks [91, 83]. However, most of these approaches are limited by their reliance on Text-Attributed Graphs (TAGs); they can only provide explanations for conditional CDGL scenarios. Extending the powerful reasoning and explanatory abilities of LLMs from TAGs to non-TAGs is also an important direction. This expansion directly affects the use of CDGL models in applications with strict security requirements, such as network fault diagnosis [9, 8].
VI-E Unified Cross-domain Evaluation Metrics.
@@
VI-F Compatibility Analysis and Improvement across Domains.
@@
VII Conclusion
The mining of commonalities in data from different domains has led CDGL to emerge as a prominent area of research in recent years, especially as it is a key point for achieving true graph foundation models. In this survey, we aim to provide an in-depth overview of existing strategies for transferring knowledge across domains. First, we introduce a novel taxonomy that categorizes techniques involving various cross-domain scales into three categories based on the different graph information transferred by specific models: structure, features, and a mixture of both. Next, we systematically review representative studies according to this taxonomy. Finally, we discuss some limitations and highlight several future research directions. Through this comprehensive review, we aspire to shed light on the advancements and challenges in the field of CDGL, thereby encouraging further progress in this domain.
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.

### v2

@@ VII Future Directions
VII-A Feature Alignment.
The correlation between feature semantics and dimensions across graph data from different domains is a crucial aspect to consider in CDGL. Misaligned feature dimensions will hinder the training of a general graph model capable of handling graph data from various domains, while misaligned semantics will prevent the trained graph model from capturing shared characteristics across domains, thereby limiting the performance of CDGL.
In existing CDGL methods, those targeting limited-scale scenarios can often skip this issue, as the graph data from source and target domains typically share identical feature dimensions by nature, such as in molecular networks [jin2024crossModalsurvey1, fang2024gpfplus]. For conditional scenarios, which mainly focus on graph data in recommendation systems [yi2023PGPRec] and TAGs [liu2024OFA, li2024zerog], feature alignment challenges are circumvented in different ways: Graph data in recommendation systems may come from different domains, but they are all built on the basis of users and items, thus avoiding feature alignment problems. For TAGs, features can be easily aligned using embeddings generated by LLMs.
However, these methods are insufficient for addressing open CDGL, where feature dimensions and semantics are highly complex and heterogeneous across domains. Although some methods have begun exploring feature alignment solutions for more open scenarios, the research is still in its early stages. For instance, methods like ProG [sun2023prog], MDGPT [yu2024MDGPT], and CDFS-GAD [chen2024CDFSGAD] have attempted to align feature dimensions across different datasets using techniques such as SVD or MLP. However, even after aligning feature dimensions, the aligned features may still lack semantic equivalence. Therefore, proposing a universal feature alignment approach to handle datasets from different domains is a key research direction towards achieving effective open CDGL.
VII-B Dealing with the Scale of Cross-domain Datasets.
@@
VII-C Overcoming data leakage.
Table I demonstrates that LLMs are receiving increasing attention in CDGL, exploring their potential capabilities in this area. However, the introduction of LLMs inevitably brings about the issue of data leakage, which has become a focal point of discussion [aiyappa2023LLMsImportant]. Specifically, since LLMs are pre-trained on extensive text corpora, it’s likely that they may have seen and memorized at least part of the test data from common benchmark datasets, especially in citation networks. This undermines the reliability of current studies that rely on earlier benchmark datasets. Moreover, [chen2024LLMsMemoryGraph] demonstrates that specific prompts could potentially enhance the ”activation” of LLMs’ corresponding memory, thereby influencing evaluation outcomes. Both [huang2023dataleakage1] and [he2024dataleakage2TAPE] have attempted to avoid the data leakage issue by collecting a new citation dataset, ensuring that the test papers are sampled from time periods after ChatGPT’s data cut-off. However, these efforts remain limited to the citation domain, and the impact of graph structures in their datasets is not significant. Therefore, it’s crucial to reconsider the methods employed to accurately evaluate the performance of LLMs on graph-related tasks. A fair, systematic, and comprehensive benchmark is also needed.
VII-D Improving interpretability.
Interpretability, also known as explainability, refers to the ability to explain or present a model’s behavior in terms that humans can understand [ijcai2024GFMSurvey]. In CDGL, most methods utilize Graph Neural Networks (GNNs). However, GNNs are typically treated as black boxes [yuan2022GNNsBlackBoxAndExplainability] and lack interpretability. While these methods consider the relationships among various pieces of information in the graph during their design and have experimentally demonstrated effectiveness in CDGL, their improvements are often based on intuition and a high-level understanding of graph data. Consequently, their interpretability remains poor. To enhance interpretability, CDGL approaches that integrate Large Language Models (LLMs) have been proposed. This is primarily due to LLMs’ reasoning and explanatory abilities, which enable them to produce user-friendly explanations for graph reasoning tasks [ijcai2024GFMSurvey, chen2024graphwiz]. However, most of these approaches are limited by their reliance on TAGs; they can only provide explanations for conditional CDGL scenarios. Extending the powerful reasoning and explanatory abilities of LLMs from TAGs to non-TAGs is also an important direction. This expansion directly affects the use of CDGL models in applications with strict security requirements, such as network fault diagnosis [zhao2023communication1, ji2022communication2].
VII-E Unified Cross-domain Evaluation Metrics.
@@
VII-F Compatibility Analysis and Improvement across Domains.
@@
VIII Conclusion
Cross-domain graph learning (CDGL) has emerged as a prominent area of research towards really general-purpose graph foundation models in recent years. In this survey, we aim to deliver an in-depth overview of current strategies that empower graph models with cross-domain transferability. Firstly, we propose a comprehensive taxonomy spanning structure-, feature-, and mixture-oriented CDGL methods across three cross-domain scales (i.e., Limited, Conditional, and Open) and three cross-domain difficulty levels (i.e., High, Moderate, and Low). Secondly, we systematically review the representative studies according to the taxonomy. Finally, we discuss some remaining limitations and pinpoint several future research directions. Through this overview, we shed light on recent advances and challenges in CDGL, and we hope to catalyse further progress in the field.
[Uncaptioned image] Haihong Zhao received his B.E. degree from The College of Software, Jilin University, Changchun, China, in 2019 and his M.sc degree from The School of Artificial Intelligence, Jilin University, Changchun, China, in 2022. He is currently a fall 2023 Ph.D. student at The Hong Kong University of Science and Technology (Guangzhou). His research interests include graph learning and data mining. He has published several papers as the leading author in top journals and conferences such as TMC, KDD, WWW, NeurIPS, and ICLR.
[Uncaptioned image] Zhixun Li received the B.Eng degree in Computer Science from Beijing Institute of Technology. He is currently a Ph.D. student in the Department of Systems Engineering and Engineering Management at the Chinese University of Hong Kong. His research interests mainly focus on deep graph learning and trustworthy AI. He has published papers at top international conferences, such as NeurIPS, SIGKDD, ICDM, and IJCAI. He has also reviewed papers in many top-tier conferences and journals, like SIGKDD, ICDM, ICDE, etc.
[Uncaptioned image] Chenyi Zi received his bachelor’s degree from Computer Science department, the South University of Technology. He is currently a fall 2023 MPhil. student at The Hong Kong University of Science and Technology (Guangzhou). His research interests include graph Pre-training, graph representation learning, and graph prompt learning.
[Uncaptioned image] Aochuan Chen received his B.E. degree from The Department of Automation, Tsinghua University, Peking, China, in 2022. He is currently a spring 2024 Ph.D. student at The Hong Kong University of Science and Technology (Guangzhou). His research interests include scalable machine learning, graph learning, and LLMs. He has published several papers as the leading author in top conferences such as CVPR, ICLR, KDD, NeurIPS, and ICML.
[Uncaptioned image] Fugee Tsung (Senior Member, IEEE) received the Ph.D. degree from the University of Michigan, Ann Arbor, MI, USA, in 1997. He is currently the Chair Professor with the Department of Industrial Engineering and Decision Analytics, The Hong Kong University of Science and Technology (HKUST), Hong Kong. He is also the Chair Professor with Data Science and Analytics Thrust and Computer Media and Arts Thrust, The Hong Kong University of Science and Technology (Guangzhou), Guangzhou, China. His research interests include quality analytics in advanced manufacturing and service processes, industrial big data, and statistical process control, monitoring, and diagnosis. Prof. Tsung is a fellow of esteemed organizations such as ASA, ASQ, IISE, IAQ, and HKIE. He is a globally recognized expert in industrial analytics and quality engineering, listed among the top 2% of most influential scientists worldwide by Stanford-Elsevier Mendeley Data in 2023 and 2024. As a Chair Professor at HKUST and HKUST (Guangzhou), he directs the Industrial and Intelligence Institute (Triple-I Institute) and the Quality and Data Analytics Laboratory (QLab). Most recently, he received the prestigious 2025 Shewhart Medal, awarded by ASQ and widely regarded as the “Nobel Prize of Quality Control.”
[Uncaptioned image] Jia Li received the PhD degree from the Chinese University of Hong Kong, in 2021. He is an assistant professor with the Hong Kong University of Science and Technology (Guangzhou) and an affiliated assistant professor with the Hong Kong University of Science and Technology. His research interests include machine learning, data mining, and deep graph learning. He has published several papers as the leading author in top conferences such as KDD, WWW and NeurIPS.
[Uncaptioned image] Jeffrey Xu YU is a Professor of Data Science and Analytics Thrust, The Hong Kong University of Science and Technology (Guangzhou). His current main research interests include graph algorithms and systems, graph mining, vector databases, query processing, and optimization. Dr. Yu has served in over 300 organization and program committees of international conferences and workshops, such as the PC Co-Chair of PAKDD’10, DASFAA’11, ICDM’12, NDBC’13, CIKM’15 and CIKM’19, and as the Conference General Co-Chair of ICDM’18. He has also served as the Information Director and Executive Committee Member of ACM SIGMOD (2007–2011), Associate Editor of IEEE Transactions on Knowledge and Data Engineering (2004–2008), Associate Editor of the VLDB Journal (2007–2013), and Chair of the Steering Committee of the Asia Pacific Web Conference (2013–2016). He is currently serving as Associate Editor for ACM Transactions on Database Systems, WWW Journal, International Journal of Cooperative Information Systems, Journal of Information Processing, and Journal on Health Information Science and Systems.